Two charges are placed on the x axis. One charge (q1 = +18.5 µC) is at x1 = +3.0 cm and the other (q2 = -11 µC) is at x2 = +9.0 cm. Find the net electric field(magnitude and direction) at the following locations.

(a) x = 0 cm
magnitude _______N/C
direction__-x direction__


(b) x = +6.0 cm
magnitude_______ N/C
direction___+x direction_
Answer #1
(a) x = 0 cm
magnitude= 136.11*10^6
direction= -x direction

(b) x = +6.0 cm

magnitude= 15*10^6
direction= + x direction
Answer #2

a) the q1 charge produces a field to the left at the origin and q2 produces one to the right) Let right be positive

so E = -k*q1/r1^2 + k*q2/r2^2 = 9.0x10^9*(18.5x10^-6/0.030^2 - 17x10^-6/0.090^2) = 1.32x10^7N/C so the magnitude is 1.321x10^7N/c and the direction is to the left

b) Now both fields produced are to the right so

E = k*q1/r1^2 + k*q2/r2^2 = 9.0x10^9*(18.5x10^-6/0.030^2 - 17x10^-6/0.030^2) = 2.75x10^8N/C to the right
Answer #3
q1 = +10.5 µC, x1 = +3.0 cm
q2 = -21 µC, x2 = +9.0 cm.
Find the net electric field (magnitude and direction) at the following locations.

(a) x = 0 cm
magnitude E = kq1/x1^2 - k|q2|/x2^2 = 8167 N/C
-x direction

(b) x = +6.0 cm
magnitude E = kq1/(x - x1)^2 + k|q2|/(x2 - x)^2 = 31500 N/C
+x direction